Solve the following system of linear equations by the substitution method. Clearly state the solution.
3x-5y=4
y=3/4x-2


Answer by Nate(3500)

y = 3/4x - 2
plug into
3x - 5y = 4
3x - 5(3/4x - 2) = 4
12/4x - 15/4x + 10 = 4
-3/4x = -6
x = 8
plug into
y = 3/4x - 2
y = 6 - 2
y = 4
(8,4)